# Shadows of the Colossus: Hierarchical Black Hole Mergers in a 10-million-body Globular Cluster Simulation

### v2

@@ 1 Introduction
Over the past decade, the LIGO/Virgo/KAGRA (LVK) network of gravitational wave (GW) detectors have been immensely successful in detecting binary black hole (BBH) mergers. Many of these GW detections have parameters that challenge typical binary star evolution scenarios, including black hole (BH) masses above the pair-instability mass gap, high spin magnitudes, and possible evidence for spin-orbit misalignment and high binary eccentricity. The recent event GW231123 (GW231123), with component masses of roughly $140\,M_{\odot}$ and $100\,M_{\odot}$ and high component spins of roughly $0.9$ and $0.8$, is a case in point. Events like GW231123 and other massive events may suggest that some of today’s GW detections may have formed via channels alternative to classic stellar evolution pathways.
@@
Recent work demonstrates the efficiency of dense stellar environments, e.g., globular clusters (GCs), young/open star clusters, and nuclear star clusters, in the dynamical formation of BBH systems (e.g., PortegiesZwartMcMillan2000; Rodriguez2016mergers; AntoniniRasio2016; Askar2017; Banerjee2017; Samsing2018; Fragione2019; DiCarlo2019; kremer2020modeling; Ye2025). These dense environments specifically enable repeated BH mergers, or hierarchical mergers, where BHs that formed in earlier BBH mergers are retained in the cluster and subsequently undergo additional mergers (e.g., MillerHamilton2002; GerosaBerti2017; Rodriguez2019; Antonini2019; FragioneRasio2023). This process provides a natural pathway for forming high-mass, high-spin BHs, and is commonly touted as a possible mechanism for forming the most massive LIGO events observed to date (e.g., Fishbach2017; Kimball2021; GW231123).
@@
The main challenge preventing the growth of BHs through hierarchical mergers is the recoil kicks attained by BH merger products, due to anisotripic emission of GWs (e.g., Bekenstein1973; Favata2004). For typical mass ratios and spin values, these kicks often exceed 100100\,km/s, and in some cases can even reach 20002000\,km/s or more (e.g., Merritt2004; Campanelli2007; Lousto2008; GerosaKesden2016). For typical GCs with masses $10^{5}-10^{6}\,M_{\odot}$ and escape velocities less than roughly 5050\,km/s (e.g., BaumgardtHilker2018), the prompt ejection of BH merger products due to these velocity kicks generally inhibit hierarchical growth beyond one, or at most two, successive merger generations. Indeed, many previous studies performing $N$-body simulations of GCs have shown that BH mergers with at least one second generation (2G) component born from a previous merger constitute at most 10% of all mergers in these systems, (e.g., Rodriguez2019).
@@
However, in the most massive GCs and nuclear star clusters with higher escape velocities, BH merger products are more easily retained, enabling significantly more high-generation hierarchical mergers (e.g., Rodriguez2020_SSC; FragioneRasio2023). For example, the Milky Way’s most massive GC, $\omega\,$Centauri, has an estimated mass of roughly $4\times 10^{6}\,M_{\odot}$ and a (present-day) central escape velocity of roughly 6060\,km/s (BaumgardtHilker2018). Haberle2024 revealed an intermediate-mass BH of mass roughly $10^{4}\,M_{\odot}$ in the center of $\omega\,$Cen, which may have formed via repeated mergers with lower-mass BHs over the cluster’s lifetime (GonzalezPrieto2025). Massive elliptical galaxies contain far more GCs relative to the Milky Way (e.g., BrodieStrader2006), including a much larger sample of very massive GCs. For example, M87 in the Virgo Supercluster is estimated to host in excess of $10^{4}$ GCs, including at least 100 GCs with masses in excess of $10^{6}\,M_{\odot}$, including a handful up to $10^{7}\,M_{\odot}$ (Jordan2009).
@@
Although the potential utility of these massive clusters in forming massive BHs via hierarchical mergers is clear, these systems pose a formidable challenge from a simulation perspective. The most massive direct N-body simulations contain up to $10^{6}$ stars, but have densities far lower than realistic systems and prohibitively long run times (e.g., wang2016dragon); however considerable recent progress has been made on this front with the PETAR code (PETAR2020; BarberAntonini2025). Even for Monte Carlo cluster methods which are considerably faster than direct N-body (e.g., Askar2017; Rodriguez2022), only a handful of massive simulations incorporating the most modern physics for BHs have been performed, often limited to high metallicities or shorter run times (for example, see the “behemoth” model of Rodriguez2020_SSC). Due to these numerical challenges, many studies have adopted analytic or semi-analytic approaches to study the BBH mergers expected in these very massive systems (e.g., AntoniniGieles2020; FragioneRasio2023; RAPSTER2024).
@@
This paper is organized as follows. In Section 2, we describe the CMC code and simulation details. In Section 3, we describe the main results from our colossus simulation, including the evolution of cluster properties and the BH population over time. In Section 4, we compare the results of colossus with lower-mass models from the CMC Cluster Catalog (kremer2020modeling), confirming several trends in cluster mass related to BH mergers and ejected binaries derived in previous literature. In Section 5, we map the observed Virgo Supercluster GCs to their closest CMC counterparts using a mass-$r_{h}$ distance function, allowing extrapolation of CMC BBH results to the wider Virgo system as a whole. We conclude and discuss in Section 6.
2 Modeling Massive Star Clusters with CMC
@@
To model our 10-million body cluster colossus, we use the Cluster Monte Carlo code CMC (Joshi2000; Pattabiraman2013; Rodriguez2022). CMC is a fully-parallelized Hénon-type (Henon1973) Monte Carlo code for stellar dynamics. This code includes all physics relevant to the evolution of dense star clusters, including two-body relaxation, tidal mass loss, stellar collisions and mergers, and direct integration of small-$N$ resonant encounters (Fregeau2007) including post-Newtonian effects (Rodriguez2018). CMC is fully coupled to stellar evolution by implementing the publicly released COSMIC software (Breivik2020), which provides rapid calculations of single and binary star evolution with up-to-date physics for massive star evolution, interacting binaries, and compact object formation. Additionally, CMC output can be converted into various observational quantities, such as surface brightness profiles, velocity-dispersion profiles, and color-magnitude diagrams, using the cmctoolkit package (Rui2021). This enables robust comparisons with realistic GCs in both the Milky Way and other galaxies. The latest catalogs of CMC models (e.g., kremer2020modeling) have proven remarkably successful in reproducing observed features of present-day GCs such as NGC 3201, NGC 6397, 47 Tuc, and others (Kremer2019a; Rui2021; Vitral2022; Ye2022). This has enabled detailed study of a range of compact object sources including radio pulsars (Ye2019), low X-ray binaries (Kremer2018a), white dwarfs (Kremer2021_wd), and binary BH mergers (Rodriguez2016mergers). For an in-depth description of the details of CMC, see Rodriguez2022.
@@
Our new $N=10^{7}$ model is intended to be an extension of the models in the CMC Cluster Catalog, therefore we adopt the same initial parameters and physical assumptions adopted there (for further detail, see kremer2020modeling). We assume an initial virial radius $r_{v}=2\,$pc, initial metallicity $Z=0.1\,Z_{\odot}$, and a galactocentric radius $R_{\rm gc}=20\,$kpc in a Milky Way-like galactic potential. All initial stellar masses are drawn from a Kroupa2001 initial mass function ranging from $0.08-150\,M_{\odot}$, yielding a total initial mass of $6.0\times 10^{6}\,M_{\odot}$. We assume an initial binary fraction of $5\%$ across all stellar mass, with binary mass ratios drawn uniformly in the range $q\in[0.1,1]$ and binary orbital periods drawn uniformly in log-space from near contact to the local hard-soft boundary. BH masses are computed using the fallback prescriptions of Fryer2012 and the (pulsational) pair-instability prescriptions of Belczynski2016. Following FullerMa2019spin, we assume all first generation BHs formed via stellar collapse are born with zero spin ($\chi=0$); we discuss this assumption further in Section 3.2. We compute spins, masses, and recoil kicks of binary BH merger products followings the method described in Rodriguez2018, which implements fits to numerical and analytic relativity calculations (Campanelli2007; Gonzalez2007; BarausseRezzolla2009; LoustoZlochower2013; GerosaKesden2016). We run the simulation for roughly 1212\,Gyr.
3 Results
3.1 Evolution of cluster parameters
@@
The half-mass radius ($r_{h}$), core radius ($r_{c}$) and Lagrange radii ($r_{L}$) of colossus are plotted over time in Figure 1. The core radius is defined using the density-weighted definition of CasertanoHut1985, which features many sharp spikes indicative of core collapse episodes of the small-$N$ black hole subcluster (e.g., BreenHeggie2013). These black hole collapses are key to the formation of black hole binaries via three-body encounters (e.g., Morscher2015), which ultimately dynamically heat the cluster as a whole (Kremer2020_bhburning). We display Lagrange radii separately for the black hole (solid lines) and non-black hole populations (dashed lines). This distinction demonstrates the decoupling of the black holes from the stars on a time scale of roughly 100100\,Myr (the mass segregation time of the black holes; e.g., Kremer2026_review), as well as the subsequent core-collapse episodes of the black hole subsystem shown in the core radius curve in the top panel.
@@
The mass, number of black holes, and number of BH binaries over time in colossus are all plotted in Figure 2, alongside five other clusters from the CMC Cluster Catalog with the same initial parameters as colossus, except for initial cluster mass. In kremer2020modeling, these are models 101, 102, 103, 104, and 146.111Note that model 146 in kremer2020modeling has metallicity $Z=0.01Z_{\odot}$, different from $Z=0.1Z_{\odot}$ adopted for colossus and models 101-104. We did not run an $N=32\times 10^{5}$ at $Z=0.1Z_{\odot}$ as part of the CMC Cluster Catalog. However, as shown in kremer2020modeling, the distinction between these two low metallicities makes little difference in the BH mass distribution nor the overall cluster evolution. The mass-loss histories of all six models (top panel of Figure 2) follow the same characteristic shape: a rapid early decline due to stellar evolutionary processes (supernovae, stellar winds, etc.), followed by a more gradual decline dominated by two-body relaxation, tidal stripping, and strong encounters in the cluster core (e.g., Weatherford2023EscapeMech). As a result, all six mass-loss curves are nearly identical in shape, differing only in a vertical offset corresponding to initial cluster mass.
@@
A comparable trend can be observed in the models’ BH population curves (middle panel of Figure 2). Since all six models utilized the same stellar initial mass function, the number of black holes throughout the end of the massive star evolution phase ($t\lesssim 10\,$Myr) scales directly with the initial cluster mass, producing identically-shaped BH population curves at early timescales. However, at later timescales where relaxation processes dominate, the BH loss rate becomes dependent on the cluster’s relaxation timescale, which scales with cluster mass (e.g., Morscher2015). Therefore, lower-mass clusters having shorter relaxation times eject their black holes more rapidly, while higher-mass clusters retain a significant BH population for much longer. colossus retained nearly two-thirds of its initial BH population by 12 Gyr, while the other models retained 1% to 50%.
@@
The number of binary black holes (bottom panel of Figure 2) reaches a maximum after roughly 1010\,Myr due to formation of BH pairs via stellar evolution of primordial binaries. We assume an initial binary fraction of $5\%$ for all models, so this initial maximum scales simply with the total number of stars. The number of BBHs remains roughly constant from $t\approx 10\,$Myr until $t\approx 100\,$Myr, when the black holes have mass segregated and formed a black hole subsystem. Once the black hole subsystem is formed, the number of BBHs decreases until reaching a value of roughly 1101-10 binaries, marking the onset of the “black hole burning” phase, in which the dynamical energy created by formation/hardening of black hole binaries roughly balances against the collapse of the cluster core (for review, see Kremer2020_bhburning). As shown, this equilibrium value scales weakly with $N$, consistent with previous studies showing that the number of BBHs in a cluster is roughly independent of the total number of BHs at late times (e.g., Chatterjee2017; Kremer2018a; MarinPinaGieles2024).
@@
3.2 Black hole mergers
@@
where $M_{\rm cl}$ is the total cluster mass and $r_{h}$ is the half-mass radius (e.g., BinneyTremaine2008). We adopt $r_{h}=4\,\text{pc}$ as a typical value for most of the cluster’s lifetime (see Figure 1), and show corresponding $v_{\rm esc}$ values on the secondary y-axis in the top panel of Figure 2. The escape velocity of colossus is near $150\,\text{km/s}$ for most of its lifetime, while all other models have $v_{\rm esc}\lesssim 100\,\text{km/s}$. This explains the discrepancy in high-generation mergers: in the lower-mass models, merger products are easily ejected by GW recoil kicks (typically $\approx 100\,\text{km/s}$), preventing hierarchical growth beyond one or two successive mergers. On the other hand, the deeper potential well of colossus and higher escape velocity threshold enable extended chains of hierarchical growth.
We note that GW recoil kick values, and therefore high-generation merger rates, are sensitive to the initial formation spins of 1G black holes. As $v_{\rm kick}$ values generally increase with the spins of merger progenitors, our assumption of $\chi_{\rm birth}=0$ for 1G black holes produces optimistic high-generation merger rates (for further discussion, see e.g., Rodriguez2019). To approximate the effect of nonzero $\chi_{\rm birth}$ on the 2G+ merger rate, we can recalculate $v_{\rm kick}$ values for all 1G mergers in each 2G merger tree assuming $\chi_{\rm birth}>0$, and consider the 2G merger suppressed if $v_{\rm kick}>v_{\rm esc}$ for any constituent merger. Assuming $\chi_{\rm birth}=0.1$ ($0.2$), we find that 44% (76%) of 2G mergers in colossus are suppressed, demonstrating the strong impact of moderate $\chi_{\rm birth}$ on the high-generation merger rate. However, the angular momentum transport prescriptions of FullerMa2019spin predict $\chi_{\rm birth}\sim 10^{-2}$ for most stellar evolution outcomes, so these suppression rates should be interpreted as upper limits. Namely, we find less than 8% 2G suppression for $\chi_{\rm birth}=0.05$, indicating our $\chi_{\rm birth}=0$ assumption remains accurate to first order.
@@
In addition, alongside each model’s BBH mergers, Figure 3 also plots the ten most massive BBH GW sources observed to date (confident events only), as reported by the Gravitational Wave Open Science Center (GWOSC). [^2] https://gwosc.org The hierarchical merger history of these observed GW events, if any, is unknown, but their locations within Figure 3 suggest a merger generation of at least two, if they are indeed hierarchical. Pulsational pair-instability mechanisms within massive stars are expected to place an upper limit of roughly $40-50~M_{\odot}$ on first-generation black holes born via massive star collapse (e.g., FowlerHoyle1964; Barkat1967; Fryer2001; Woosley2017). In the CMC models described here, we assume stars that undergo pulsational-pair instability supernovae yield black holes of mass precisely $40.5\,M_{\odot}$, following the prescriptions of Belczynski2016. In this case, ignoring contributions from major stellar collisions (Kremer2020), the highest mass a generation 2 black hole can have in a CMC model is approximately $80~M_{\odot}$. This limit can be observed in Figure 3 as a primary mass cutoff of about $80~M_{\odot}$ for generation 2 mergers. Since three of the most massive observed GW events plotted in Figure 3 have a primary mass exceeding $80~M_{\odot}$, these mergers would need to be generation 3 or higher (assuming stellar collisions or accretion do not contribute significantly, see Kiroglu2025c, for further discussion). Since generation 3 mergers are demonstrably rare in low-mass GCs, these observed GW events are much more likely to come from a GC with mass similar to colossus, rather than a typical lower-mass GC.

### v3

@@ I Introduction
Over the past decade, the LIGO/Virgo/KAGRA (LVK) network of gravitational wave (GW) detectors have been immensely successful in detecting binary black hole (BBH) mergers. Many of these GW detections have parameters that challenge typical binary star evolution scenarios, including black hole (BH) masses above the pair-instability mass gap, high spin magnitudes, and possible evidence for spin-orbit misalignment and high binary eccentricity. The recent event GW231123 (The LIGO Scientific Collaboration et al., 2025), with component masses of roughly $140\,M_{\odot}$ and $100\,M_{\odot}$ and high component spins of roughly $0.9$ and $0.8$, is a case in point. Events like GW231123 and other massive events may suggest that some of today’s GW detections may have formed via channels alternative to classic stellar evolution pathways.
@@
Recent work demonstrates the efficiency of dense stellar environments, e.g., globular clusters (GCs), young/open star clusters, and nuclear star clusters, in the dynamical formation of BBH systems (e.g., Portegies Zwart and McMillan, 2000; Rodriguez et al., 2016; Antonini and Rasio, 2016; Askar et al., 2017; Banerjee, 2017; Samsing, 2018; Fragione et al., 2019; Di Carlo and others, 2019; Kremer et al., 2020b; Ye et al., 2025). These dense environments specifically enable repeated BH mergers, or hierarchical mergers, where BHs that formed in earlier BBH mergers are retained in the cluster and subsequently undergo additional mergers (e.g., Miller and Hamilton, 2002; Gerosa and Berti, 2017; Rodriguez et al., 2019; Antonini et al., 2019; Fragione and Rasio, 2023). This process provides a natural pathway for forming high-mass, high-spin BHs, and is commonly touted as a possible mechanism for forming the most massive LIGO events observed to date (e.g., Fishbach et al., 2017; Kimball et al., 2021; The LIGO Scientific Collaboration et al., 2025).
@@
The main challenge preventing the growth of BHs through hierarchical mergers is the recoil kicks attained by BH merger products, due to anisotripic emission of GWs (e.g., Bekenstein, 1973; Favata et al., 2004). For typical mass ratios and spin values, these kicks often exceed 100100\,km/s, and in some cases can even reach 20002000\,km/s or more (e.g., Merritt et al., 2004; Campanelli et al., 2007; Lousto and Zlochower, 2008; Gerosa and Kesden, 2016). For typical GCs with masses $10^{5}-10^{6}\,M_{\odot}$ and escape velocities less than roughly 5050\,km/s (e.g., Baumgardt and Hilker, 2018), the prompt ejection of BH merger products due to these velocity kicks generally inhibit hierarchical growth beyond one, or at most two, successive merger generations. Indeed, many previous studies performing $N$-body simulations of GCs have shown that BH mergers with at least one second generation (2G) component born from a previous merger constitute at most 10% of all mergers in these systems, (e.g., Rodriguez et al., 2019).
@@
However, in the most massive GCs and nuclear star clusters with higher escape velocities, BH merger products are more easily retained, enabling significantly more high-generation hierarchical mergers (e.g., Rodriguez et al., 2020; Fragione and Rasio, 2023). For example, the Milky Way’s most massive GC, $\omega\,$Centauri, has an estimated mass of roughly $4\times 10^{6}\,M_{\odot}$ and a (present-day) central escape velocity of roughly 6060\,km/s (Baumgardt and Hilker, 2018). Häberle et al. (2024) revealed an intermediate-mass BH of mass roughly $10^{4}\,M_{\odot}$ in the center of $\omega\,$Cen, which may have formed via repeated mergers with lower-mass BHs over the cluster’s lifetime (González Prieto et al., 2025). Massive elliptical galaxies contain far more GCs relative to the Milky Way (e.g., Brodie and Strader, 2006), including a much larger sample of very massive GCs. For example, M87 in the Virgo Supercluster is estimated to host in excess of $10^{4}$ GCs, including at least 100 GCs with masses in excess of $10^{6}\,M_{\odot}$, including a handful up to $10^{7}\,M_{\odot}$ (Jordán et al., 2009).
@@
Although the potential utility of these massive clusters in forming massive BHs via hierarchical mergers is clear, these systems pose a formidable challenge from a simulation perspective. The most massive direct N-body simulations contain up to $10^{6}$ stars, but have densities far lower than realistic systems and prohibitively long run times (e.g., Wang et al., 2016); however considerable recent progress has been made on this front with the PETAR code (Wang et al., 2020; Barber and Antonini, 2025). Even for Monte Carlo cluster methods which are considerably faster than direct N-body (e.g., Askar et al., 2017; Rodriguez and others, 2022), only a handful of massive simulations incorporating the most modern physics for BHs have been performed, often limited to high metallicities or shorter run times (for example, see the “behemoth” model of Rodriguez et al., 2020). Due to these numerical challenges, many studies have adopted analytic or semi-analytic approaches to study the BBH mergers expected in these very massive systems (e.g., Antonini and Gieles, 2020; Fragione and Rasio, 2023; Kritos et al., 2024).
@@
This paper is organized as follows. In Section II, we describe the CMC code and simulation details. In Section III, we describe the main results from our colossus simulation, including the evolution of cluster properties and the BH population over time. In Section IV, we compare the results of colossus with lower-mass models from the CMC Cluster Catalog (Kremer et al., 2020b), confirming several trends in cluster mass related to BH mergers and ejected binaries derived in previous literature. In Section V, we map the observed Virgo Supercluster GCs to their closest CMC counterparts using a mass-$r_{h}$ distance function, allowing extrapolation of CMC BBH results to the wider Virgo system as a whole. We conclude and discuss in Section VI.
II Modeling Massive Star Clusters with CMC
@@
To model our 10-million body cluster colossus, we use the Cluster Monte Carlo code CMC (Joshi et al., 2000; Pattabiraman et al., 2013; Rodriguez and others, 2022). CMC is a fully-parallelized Hénon-type (Henon, 1973) Monte Carlo code for stellar dynamics. This code includes all physics relevant to the evolution of dense star clusters, including two-body relaxation, tidal mass loss, stellar collisions and mergers, and direct integration of small-$N$ resonant encounters (Fregeau and Rasio, 2007) including post-Newtonian effects (Rodriguez et al., 2018). CMC is fully coupled to stellar evolution by implementing the publicly released COSMIC software (Breivik et al., 2020), which provides rapid calculations of single and binary star evolution with up-to-date physics for massive star evolution, interacting binaries, and compact object formation. Additionally, CMC output can be converted into various observational quantities, such as surface brightness profiles, velocity-dispersion profiles, and color-magnitude diagrams, using the cmctoolkit package (Rui et al., 2021). This enables robust comparisons with realistic GCs in both the Milky Way and other galaxies. The latest catalogs of CMC models (e.g., Kremer et al., 2020b) have proven remarkably successful in reproducing observed features of present-day GCs such as NGC 3201, NGC 6397, 47 Tuc, and others (Kremer et al., 2019a; Rui et al., 2021; Vitral et al., 2022; Ye et al., 2022). This has enabled detailed study of a range of compact object sources including radio pulsars (Ye et al., 2019), low X-ray binaries (Kremer et al., 2018), white dwarfs (Kremer et al., 2021), and binary BH mergers (Rodriguez et al., 2016). For an in-depth description of the details of CMC, see Rodriguez and others (2022).
@@
Our new $N=10^{7}$ model is intended to be an extension of the models in the CMC Cluster Catalog, therefore we adopt the same initial parameters and physical assumptions adopted there (for further detail, see Kremer et al., 2020b). We assume an initial virial radius $r_{v}=2\,$pc, initial metallicity $Z=0.1\,Z_{\odot}$, and a galactocentric radius $R_{\rm gc}=20\,$kpc in a Milky Way-like galactic potential. All initial stellar masses are drawn from a Kroupa (2001) initial mass function ranging from $0.08-150\,M_{\odot}$, yielding a total initial mass of $6.0\times 10^{6}\,M_{\odot}$. We assume an initial binary fraction of $5\%$ across all stellar mass, with binary mass ratios drawn uniformly in the range $q\in[0.1,1]$ and binary orbital periods drawn uniformly in log-space from near contact to the local hard-soft boundary. BH masses are computed using the fallback prescriptions of Fryer et al. (2012) and the (pulsational) pair-instability prescriptions of Belczynski and others (2016). Following Fuller and Ma (2019), we assume all first generation BHs formed via stellar collapse are born with zero spin ($\chi=0$); we discuss this assumption further in Section III.2. We compute spins, masses, and recoil kicks of binary BH merger products followings the method described in Rodriguez et al. (2018), which implements fits to numerical and analytic relativity calculations (Campanelli et al., 2007; González et al., 2007; Barausse and Rezzolla, 2009; Lousto and Zlochower, 2007; Gerosa and Kesden, 2016). We run the simulation for roughly 1212\,Gyr.
III Results
III.1 Evolution of cluster parameters
@@
The half-mass radius ($r_{h}$), core radius ($r_{c}$) and Lagrange radii ($r_{L}$) of colossus are plotted over time in Figure 1. The core radius is defined using the density-weighted definition of Casertano and Hut (1985), which features many sharp spikes indicative of core collapse episodes of the small-$N$ black hole subcluster (e.g., Breen and Heggie, 2013). These black hole collapses are key to the formation of black hole binaries via three-body encounters (e.g., Morscher and others, 2015), which ultimately dynamically heat the cluster as a whole (Kremer et al., 2020a). We display Lagrange radii separately for the black hole (solid lines) and non-black hole populations (dashed lines). This distinction demonstrates the decoupling of the black holes from the stars on a time scale of roughly 100100\,Myr (the mass segregation time of the black holes; e.g., Kremer, 2026), as well as the subsequent core-collapse episodes of the black hole subsystem shown in the core radius curve in the top panel.
@@
The mass, number of black holes, and number of BH binaries over time in colossus are all plotted in Figure 2, alongside five other clusters from the CMC Cluster Catalog with the same initial parameters as colossus, except for initial cluster mass. In Kremer et al. (2020b), these are models 101, 102, 103, 104, and 146.111Note that model 146 in Kremer et al. (2020b) has metallicity $Z=0.01Z_{\odot}$, different from $Z=0.1Z_{\odot}$ adopted for colossus and models 101-104. We did not run an $N=32\times 10^{5}$ at $Z=0.1Z_{\odot}$ as part of the CMC Cluster Catalog. However, as shown in Kremer et al. (2020b), the distinction between these two low metallicities makes little difference in the BH mass distribution nor the overall cluster evolution. The mass-loss histories of all six models (top panel of Figure 2) follow the same characteristic shape: a rapid early decline due to stellar evolutionary processes (supernovae, stellar winds, etc.), followed by a more gradual decline dominated by two-body relaxation, tidal stripping, and strong encounters in the cluster core (e.g., Weatherford et al., 2023). As a result, all six mass-loss curves are nearly identical in shape, differing only in a vertical offset corresponding to initial cluster mass.
@@
A comparable trend can be observed in the models’ BH population curves (middle panel of Figure 2). Since all six models utilized the same stellar initial mass function, the number of black holes throughout the end of the massive star evolution phase ($t\lesssim 10\,$Myr) scales directly with the initial cluster mass, producing identically-shaped BH population curves at early timescales. However, at later timescales where relaxation processes dominate, the BH loss rate becomes dependent on the cluster’s relaxation timescale, which scales with cluster mass (e.g., Morscher and others, 2015). Therefore, lower-mass clusters having shorter relaxation times eject their black holes more rapidly, while higher-mass clusters retain a significant BH population for much longer. colossus retained nearly two-thirds of its initial BH population by 12 Gyr, while the other models retained 1% to 50%.
@@
The number of binary black holes (bottom panel of Figure 2) reaches a maximum after roughly 1010\,Myr due to formation of BH pairs via stellar evolution of primordial binaries. We assume an initial binary fraction of $5\%$ for all models, so this initial maximum scales simply with the total number of stars. The number of BBHs remains roughly constant from $t\approx 10\,$Myr until $t\approx 100\,$Myr, when the black holes have mass segregated and formed a black hole subsystem. Once the black hole subsystem is formed, the number of BBHs decreases until reaching a value of roughly 1101-10 binaries, marking the onset of the “black hole burning” phase, in which the dynamical energy created by formation/hardening of black hole binaries roughly balances against the collapse of the cluster core (for review, see Kremer et al., 2020a). As shown, this equilibrium value scales weakly with $N$, consistent with previous studies showing that the number of BBHs in a cluster is roughly independent of the total number of BHs at late times (e.g., Chatterjee et al., 2017; Kremer et al., 2018; Marín Pina and Gieles, 2024).
@@
III.2 Black hole mergers
@@
where $M_{\rm cl}$ is the total cluster mass and $r_{h}$ is the half-mass radius (e.g., Binney and Tremaine, 2008). We adopt $r_{h}=4\,\text{pc}$ as a typical value for most of the cluster’s lifetime (see Figure 1), and show corresponding $v_{\rm esc}$ values on the secondary y-axis in the top panel of Figure 2. The escape velocity of colossus is near $150\,\text{km/s}$ for most of its lifetime, while all other models have $v_{\rm esc}\lesssim 100\,\text{km/s}$. This explains the discrepancy in high-generation mergers: in the lower-mass models, merger products are easily ejected by GW recoil kicks (typically $\approx 100\,\text{km/s}$), preventing hierarchical growth beyond one or two successive mergers. On the other hand, the deeper potential well of colossus and higher escape velocity threshold enable extended chains of hierarchical growth.
We note that GW recoil kick values, and therefore high-generation merger rates, are sensitive to the initial formation spins of 1G black holes. As $v_{\rm kick}$ values generally increase with the spins of merger progenitors, our assumption of $\chi_{\rm birth}=0$ for 1G black holes produces optimistic high-generation merger rates (for further discussion, see e.g., Rodriguez et al., 2019). To approximate the effect of nonzero $\chi_{\rm birth}$ on the 2G+ merger rate, we can recalculate $v_{\rm kick}$ values for all 1G mergers in each 2G merger tree assuming $\chi_{\rm birth}>0$, and consider the 2G merger suppressed if $v_{\rm kick}>v_{\rm esc}$ for any constituent merger. Assuming $\chi_{\rm birth}=0.1$ ($0.2$), we find that 44% (76%) of 2G mergers in colossus are suppressed, demonstrating the strong impact of moderate $\chi_{\rm birth}$ on the high-generation merger rate. However, the angular momentum transport prescriptions of Fuller and Ma (2019) predict $\chi_{\rm birth}\sim 10^{-2}$ for most stellar evolution outcomes, so these suppression rates should be interpreted as upper limits. Namely, we find less than 8% 2G suppression for $\chi_{\rm birth}=0.05$, indicating our $\chi_{\rm birth}=0$ assumption remains accurate to first order.
@@
In addition, alongside each model’s BBH mergers, Figure 3 also plots the ten most massive BBH GW sources observed to date (confident events only), as reported by the Gravitational Wave Open Science Center (GWOSC). [^2] https://gwosc.org The hierarchical merger history of these observed GW events, if any, is unknown, but their locations within Figure 3 suggest a merger generation of at least two, if they are indeed hierarchical. Pulsational pair-instability mechanisms within massive stars are expected to place an upper limit of roughly $40-50~M_{\odot}$ on first-generation black holes born via massive star collapse (e.g., Fowler and Hoyle, 1964; Barkat et al., 1967; Fryer et al., 2001; Woosley, 2017). In the CMC models described here, we assume stars that undergo pulsational-pair instability supernovae yield black holes of mass precisely $40.5\,M_{\odot}$, following the prescriptions of Belczynski and others (2016). In this case, ignoring contributions from major stellar collisions (Kremer and others, 2020), the highest mass a generation 2 black hole can have in a CMC model is approximately $80~M_{\odot}$. This limit can be observed in Figure 3 as a primary mass cutoff of about $80~M_{\odot}$ for generation 2 mergers. Since three of the most massive observed GW events plotted in Figure 3 have a primary mass exceeding $80~M_{\odot}$, these mergers would need to be generation 3 or higher (assuming stellar collisions or accretion do not contribute significantly, see Kıroğlu et al., 2025b, for further discussion). Since generation 3 mergers are demonstrably rare in low-mass GCs, these observed GW events are much more likely to come from a GC with mass similar to colossus, rather than a typical lower-mass GC.